translate to an algebraic equation then solve. the lenght of a rectangle is six units longer than the width. find the lenght and width if the perimeter of the rectangle is 60 units.
2) fin the circumference and area of a circle with the radius 3.5 feet. use 22/7 for pi.
3) the width of a rectangular playground is 18 meters less than the lenght. find the width when the perimeter of the rectangle is 84 meters.
once again thanks a lot this will help me get my ged

Answer by stanbon(75887)

translate to an algebraic equation then solve. the length of a rectangle is six units longer than the width. find the length and width if the perimeter of the rectangle is 60 units.
Let width = x units
Then length = x+6 units
Perimeter = 2(length + width)
2(x+(x+6)) = 60
2x+6 = 30
2x = 24
x = 12 units (width)
x+6 = 18 units (length)
-----------------------
2) find the circumference and area of a circle with the radius 3.5 feet. use 22/7 for pi.
Area = (pi)r^2
Area = (22/7)*3.5^2 = 22*3.5*3.5/7 = 22*3.5/2 = 11*3.5 = 38.5 sq ft
-----------------------
3) the width of a rectangular playground is 18 meters less than the lenght. find the width when the perimeter of the rectangle is 84 meters.
---------------
Let the length be "x meters; the width is "x-18" meters.
Perimeter = 2(length + width)
84 = 2(x + x-18)
2x-18 = 42
2x = 60
x = 30 meters (length)
x-18 = 12 meters (width)
================
Cheers,
Stan H.